OpenStudy (anonymous):

Need help with finding the hyperplane for a=<-2, 1 ,4>

OpenStudy (anonymous):

like for a=<0, -3, 5, 7> I don't know why the answer is (1,0,0,0),(7,5,3,0),(0,7,0,3)

OpenStudy (zarkon):

for your example isn't the answer the span of those 3 vectors

OpenStudy (anonymous):

Okay how would you get the span of these 3 vectors then so frustrated I tried this so many times already

OpenStudy (zarkon):

Am I also assume you are trying to find the hyperplane that goes through the origin?

OpenStudy (anonymous):

yes

OpenStudy (zarkon):

typically the hyperplane through the point P is the set of all vectors y such that \[n\cdot(y-p)=0\] where n is called the normal vector

OpenStudy (zarkon):

so you need to find all the vectors y such that \[a\cdot y=0\]

OpenStudy (zarkon):

have you done the Gram Schmidt orthoginalization algorithm before?

OpenStudy (anonymous):

Okay this is what I did I maade -2x_1 + x_2 + 4x_3 = 0

OpenStudy (anonymous):

(-2x_1 + x_2 + 4x_3, x_2,x_3)=x_1(-2,0,0) + x_2(1,1,0) + x_3(4,0,1)

OpenStudy (anonymous):

then v_1 =(-2 0 0 ) v_2=(1 1 0) v_3 =(4 0 1)

OpenStudy (anonymous):

is that wrong? it looks wrong for some reason

OpenStudy (zarkon):

Sorry I have to go.... it does look wrong...look into the algorithm I mentioned

OpenStudy (zarkon):

try this ..pick any vector that is orthogonal to <-2, 1 ,4> like <1,2,0> now pick another that is orthogonal to <-2, 1 ,4> <2,0,1> <1,2,0> and <2,0,1> are clearly independent of each other so the hyperplane is the span(<1,2,0>,<2,0,1>)